Help

Help
media%2F28c%2F28ccfd9c-a6b5-4d1a-bff6-a0
Answer #1

answer1) a)we know the formula

m=\rhoV=1060kg/m3*6.2*10-3m3=6.57 kg

and we know weight

W=mg=6.57*9.8m/s2=64N

b) and the percentage=(64/750)*100=8.5%

answer2)

here we have \DeltaP=\rhogh=1060kg/m3*9.8m/s2*1.43=1.485*104Pa

so answer is 1.485*104Pa

answer3)

a) we know the formula

Q=Av, Q=V/t=10*10-3/31=0.32258*10-3

so v=Q/A=0.32258*10-3/3.1*10-4=1.04m/s

b) from continuity eqn we know

A1v1=A2v2

v2=A1v1/A2..........1)

here A2=A1/2

so eqn 1 becomes

v2=2v1 ( A1 cancels out)

v2=2*1.041=2.08 m/s